# TempFlow-GRPO: When Timing Matters for GRPO in Flow Models

### v2

@@ Abstract
Recent flow matching models for text-to-image generation have achieved remarkable quality, yet their integration with reinforcement learning for human preference alignment remains suboptimal, hindering fine-grained reward-based optimization. We observe that the key impediment to effective GRPO training of flow models is the temporal uniformity assumption in existing approaches: sparse terminal rewards with uniform credit assignment fail to capture the varying criticality of decisions across generation timesteps, resulting in inefficient exploration and suboptimal convergence. To remedy this shortcoming, we introduce TempFlow-GRPO (Temporal Flow GRPO), a principled GRPO framework that captures and exploits the temporal structure inherent in flow-based generation. TempFlow-GRPO introduces three key innovations: (i) a trajectory branching mechanism that provides process rewards by concentrating stochasticity at designated branching points, enabling precise credit assignment without requiring specialized intermediate reward models; (ii) a noise-aware weighting scheme that modulates policy optimization according to the intrinsic exploration potential of each timestep, prioritizing learning during high-impact early stages while ensuring stable refinement in later phases; and (iii) a seed group strategy that controls for initialization effects to isolate exploration contributions. These innovations endow the model with temporally-aware optimization that respects the underlying generative dynamics, leading to state-of-the-art performance in human preference alignment and text-to-image benchmarks.
@@ 1 Introduction
While text-to-image diffusion models have achieved unprecedented visual quality and semantic control Esser et al. (2024); Xie et al. (2024a); Labs et al. (2025), aligning their outputs with human preference remains a formidable challenge. Reinforcement learning has emerged as a promising solution, giving rise to the field of Diffusion RL Wallace et al. (2024); Black et al. (2023); Fan et al. (2023). However, the performance of these methods remains suboptimal, hindered by two fundamental limitations that have been largely overlooked: ignoring the temporal dynamics of generation and lacking intermediate feedback signals. These approaches apply uniform optimization across all timesteps and provide rewards only at completion, missing the varying importance of decisions throughout the generation process.
The majority of existing approaches Gu et al. (2024); Hong et al. (2024), including recent works like Flow-GRPO Liu et al. (2025) and DanceGRPO Xue et al. (2025), treat the multi-step generation process as a ”black box” with temporally agnostic optimization. They apply uniform updates across all timesteps despite the fact that each timestep operates under different noise conditions and contributes differently to final image quality. Specifically, we plot the Figure 2 (left) with applying SDE at only one timestep in the entire ODE trajectory, which ensures that any deviation in the final reward can be attributed to the stochastic exploration introduced at that specific step. As shown in Figure 2 (left), the std of the reward varies dramatically between timesteps, reaching a peak during early structural decisions (steps 0-2) and approaching zero during final refinements (steps 6-8). Yet Flow-GRPO maintains uniform treatment throughout, squandering high-impact exploration opportunities. Alternative approaches like SPO Liang et al. (2025) attempt to address temporal dynamics through process reward models, but training such models on semantically ambiguous intermediate states is notoriously difficult. This raises a fundamental question: how can we effectively achieve precise credit assignment for intermediate actions while adapting optimization intensity to each timestep’s exploration capacity?
@@ 2 Related Work
Alignment for Diffusion Models. Alignment for diffusion models has emerged as a key research area. D3PO Yang et al. (2024) introduces Direct Preference for Denoising Diffusion Policy Optimization to directly fine-tune diffusion models. Diffusion-DPO Wallace et al. (2024) adapts Direct Preference Optimization (DPO Rafailov et al. (2023)) as a simpler RLHF alternative that directly optimizes policies satisfying human preferences. DyMO Xie & Gong (2025) proposes a training-free alignment method for inference-time preference alignment. Flow-GRPO Liu et al. (2025) first integrates online reinforcement learning into flow matching models. However, Flow-GRPO applies uniform optimization across timesteps and suffers from sparse terminal rewards, ignoring time-varying exploration potential in stochastic processes. Our TempFlow-GRPO addresses these limitations through trajectory branching for precise credit assignment and noise-aware policy weighting aligned with natural exploration capacity.
@@
Process Reward. Shaping reward processes beyond sparse terminal rewards significantly improves learning and policy performance. Zhang et al. Zhang et al. (2025) combines response-level and step-level evaluation metrics. ThinkPRM Khalifa et al. (2025) builds data-efficient process reward models (PRMs) using verification chain-of-thought for step-wise verification. In diffusion models, SPO Liang et al. (2024) trains step-aware preference models for both noisy and clean images. However, PRMs require expensive step-level supervision. Methods like PRIME Cui et al. (2025) enable online PRM updates using only outcome labels through implicit process rewards. Math-Shepherd Wang et al. (2023) assigns reward scores to solution steps. Given challenges in scoring noisy images, algorithms enabling process rewards for flow models are needed. Our TempFlow-GRPO circumvents specialized process reward models by directly attributing outcome-based signals to intermediate actions, enabling precise credit assignment without computational overhead of training step-level evaluators for semantically ambiguous states.
@@ 3 Preliminary: Flow-GRPO
Given a prompt $\bm{c}$, the flow model $p_{\theta}$ samples a group of $G$ individual images $\{\bm{x}_{0}^{i}\}_{i=1}^{G}$ and the corresponding reverse-time trajectories $\{(\bm{x}_{T}^{i},\bm{x}_{T-1}^{i},...,\bm{x}_{0}^{i}\}_{i=1}^{G}$. Then, the advantage of the $i$-th image is calculated by normalizing the group-level rewards as follows:
@@
Figure 5: Overview of TempFlow-GRPO Framework. Our method performs trajectory branching by switching from ODE to SDE sampling at selected timesteps (t=k, j, i), injecting noise $\sigma_{t}\sqrt{\Delta t}\epsilon$ to create exploratory branches. Each branch generates a distinct outcome with reward $R_{i}$, enabling precise credit assignment. The framework applies noise-aware weighting where $\omega_{i}>\omega_{j}>\omega_{k}$, prioritizing optimization at high-noise early stages (larger circles) over low-noise refinement phases (smaller circles), aligning learning intensity with each timestep’s intrinsic exploration capacity. We visualize the model’s learning process as an astronaut exploring unknown planets: in early stages , the model explores vast possibility spaces with high uncertainty, while later stages involve focused navigation toward the final destination.
@@ 4.1 Temporal Flow-GRPO
Flow-GRPO advances online RL for flow matching but overlooks time-dependent generative dynamics. We identify two key limitations: (1) Sparse Terminal Rewards: uniform credit assignment across timesteps fails to distinguish critical early decisions from later fine-tuning; (2) Uniform Optimization: ignoring non-uniform noise in SDE sampling, where high-noise early steps offer greater exploration potential than low-noise refinement phases Xie & Gong (2025). To address these issues, we propose TempFlow-GRPO (Figure 5), introducing process rewards via trajectory branching and noise-aware policy weighting.
@@ 4.1.2 Noise-Aware Policy Weighting
Reweighting by Noise Level. We wonder if directly leverage the level of the exploration space itself as a proxy for a reweighting factor. We visualized the noise level alongside the reward standard deviation (Figure 6, left) and observed a striking correspondence between the two. This strong correlation suggests that the noise level, serves as an excellent and intrinsic proxy for the exploration capacity and associated risk at each timestep.
We therefore propose to reweight the policy loss directly using the noise level. For each timestep $t$, we introduce a weighting factor proportional to the noise level. Specifically, we modify the original policy gradient loss function to the following weighted form:
@@ 4.3 Group Strategy
Batch Group & Prompt Group. The original GRPO Shao et al. (2024) framework employs a group-level strategy, where trajectories are grouped based on a shared prompt. To mitigate issues of overfitting and instability, Reinforce++ Hu et al. (2025) later introduced a batch-level normalization.
Seed Group. In our work, trajectory branching requires $K$ distinct explorations at each timestep. Furthermore, we propose a novel seed-level grouping strategy. Under this approach, trajectories originating from the same prompt are further grouped if they share an identical initial noise seed. This methodology effectively controls for the influence of the initial noise, thereby ensuring that variations in reward can be attributed solely to the exploration conducted during the branching process. The experimental results, presented in Figure 7, validate our approach. As shown, TempFlow-GRPO consistently achieves superior performance compared to Flow-GRPO, regardless of the grouping strategy employed.
@@ 5 Experiment
Following Flow-GRPO, we validate our approach on the Compositional Image Generation benchmark in Geneval Ghosh et al. (2023) and the Human Preference Alignment task in PickScore Kirstain et al. (2023). To ensure a fair comparison, our experimental setup is kept consistent with that of Flow-GRPO. For instance, we normalized the weights applied to the policy loss to have a mean of 1 at all timesteps. While our method introduces a trajectory branching strategy, we maintain an identical group configuration by setting 4 initial noise seeds and a branching factor ($K$) of 6. This results in a total group size of 24 (4 seeds $\times$ 6 branches) and 48 groups, perfectly matching the setup in Flow-GRPO. Furthermore, to benchmark our method against concurrent work, we conduct additional experiments on the FLUX.1-dev model. More detailed experimental settings (include comparsion with DanceGRPO) can be found in the Appendix A.2.
@@ 5.1 Main Results
LDM Rombach et al. (2022) - 0.37 0.92 0.29 0.23 0.70 0.02 0.05
SD1.5 Rombach et al. (2022) - 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD2.1 Rombach et al. (2022) - 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL Podell et al. (2023) - 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALLE-2 Ramesh et al. (2022) - 0.52 0.94 0.66 0.49 0.77 0.10 0.19
DALLE-3 Betker et al. (2023) - 0.67 0.96 0.87 0.47 0.83 0.43 0.45
@@
Show-o Xie et al. (2024b) - 0.53 0.95 0.52 0.49 0.82 0.11 0.28
Emu3-Gen Wang et al. (2024) - 0.54 0.98 0.71 0.34 0.81 0.17 0.21
JanusFlow Ma et al. (2025a) - 0.63 0.97 0.59 0.45 0.83 0.53 0.42
Janus-Pro-7B Chen et al. (2025) - 0.80 0.99 0.89 0.59 0.90 0.79 0.66
GPT-4o Hurst et al. (2024) - 0.84 0.99 0.92 0.85 0.92 0.75 0.61
@@
FLUX.1 Dev Black et al. (2025) - 0.66 0.98 0.81 0.74 0.79 0.22 0.45
SD3.5-L Esser et al. (2024) - 0.71 0.98 0.89 0.73 0.83 0.34 0.47
SANA-1.5 4.8B Xie et al. (2025) - 0.81 0.99 0.93 0.86 0.84 0.59 0.65
SD3.5-M Esser et al. (2024) - 0.63 0.98 0.78 0.50 0.81 0.24 0.52
@@
SD3.5-M+Flow-GRPO Liu et al. (2025) 5600 0.95 1.00 0.99 0.95 0.92 0.99 0.86
SD3.5-M+Flow-GRPO Liu et al. (2025) 3800 0.88 0.99 0.96 0.90 0.87 0.83 0.78
@@
HPSv3 on FLUX.1-dev. To further validate the performance of TempFlow-GRPO across different model and reward models, we select FLUX.1-dev as our base model with 1024 resolution and employ HPSv3 as the reward model. As shown in Figure 8, which presents the results on PickScore, our method requires only 80 steps to match the performance that Flow-GPO achieves in 300 steps. Moreover, our approach maintains a lower and more stable KL loss. Further experimental results and visualizations for the FLUX model can be found in the Appendix. A.3.
@@ 5.2 Analysis
Qualitative Result. We also conducted qualitative analyses on the FLUX.1-dev, Flow-GRPO (Prompt), and TempFlow-GRPO. As shown in Figure10, compared to Flow-GRPO (Prompt), TempFlow-GRPO produces images with noticeably finer details and fewer visual artifacts or mistakes. In particular, our approach demonstrates superior capability in preserving complex structures and realistic textures. These qualitative improvements further highlighting the advantages of our method in generating high-quality, visually appealing images.
@@ 6 Conclusion
We utilize LLMs to assist with formula derivations and writing refinement on this paper.
@@ A.1 Policy Gradient-Based Theoretical Framework
This reveals that the natural gradient coefficient is proportional to $\sqrt{\frac{1-k}{k}}\sqrt{\Delta k}$, which captures the intrinsic exploration potential at timestep $k$. After reweighting, we have the following derivation:
@@ A.2 Experimental Setting Details
Dataset. We evaluate Compositional Image Generation on Geneval and Human Preference Alignment on PickScore. Meanwhile, we also test our method on HPDv2 Wu et al. (2023).
Reward Model. In TempFlow-GRPO, we introduce four reward models, including clip-based PickScore, HPSv2 Wu et al. (2023), VLM-based HPSv3 Ma et al. (2025b) and framework Geneval.
Compositional Image Generation. Following Flow-GRPO, we evaluate our method on PickScore prompts using the PickScore reward function. PickScore is a CLIP-based scoring function that demonstrates superhuman performance in predicting human preferences for generated images. To ensure consistency with Flow-GRPO, we set the KL divergence regularization coefficient $\beta=0.001$.
Human Preference Alignment. We adopt the GenEval framework following Flow-GRPO’s experimental protocol. The training prompts are sourced from the Flow-GRPO dataset. GenEval provides an object-focused evaluation framework that assesses compositional image properties, including object co-occurrence, spatial positioning, object count, and color attributes. For fair comparison, we maintain the KL divergence regularization coefficient at $\beta=0.004$.

### v3

@@ Abstract
Recent flow matching models for text-to-image generation have achieved remarkable quality, yet their integration with reinforcement learning for human preference alignment remains suboptimal, hindering fine-grained reward-based optimization. We observe that the key impediment to effective GRPO training of flow models is the temporal uniformity assumption in existing approaches: sparse terminal rewards with uniform credit assignment fail to capture the varying criticality of decisions across generation timesteps, resulting in inefficient exploration and suboptimal convergence. To remedy this shortcoming, we introduce TempFlow-GRPO (Temporal Flow-GRPO), a principled GRPO framework that captures and exploits the temporal structure inherent in flow-based generation. TempFlow-GRPO introduces three key innovations: (i) a trajectory branching mechanism that provides process rewards by concentrating stochasticity at designated branching points, enabling precise credit assignment without requiring specialized intermediate reward models; (ii) a noise-aware weighting scheme that modulates policy optimization according to the intrinsic exploration potential of each timestep, prioritizing learning during high-impact early stages while ensuring stable refinement in later phases; and (iii) a seed group strategy that controls for initialization effects to isolate exploration contributions. These innovations endow the model with temporally-aware optimization that respects the underlying generative dynamics, leading to state-of-the-art performance in human preference alignment and text-to-image benchmarks.
@@ 1 Introduction
While text-to-image diffusion models have achieved unprecedented visual quality and semantic control (Esser et al., 2024; Xie et al., 2024a; Labs et al., 2025), aligning their outputs with human preference remains a formidable challenge. Reinforcement learning has emerged as a promising solution, giving rise to the field of Diffusion RL (Wallace et al., 2024; Black et al., 2023; Fan et al., 2023). However, the performance of these methods remains suboptimal, hindered by two fundamental limitations that have been largely overlooked: ignoring the temporal dynamics of generation and lacking intermediate feedback signals. These approaches apply uniform optimization across all timesteps and provide rewards only at completion, missing the varying importance of decisions throughout the generation process.
The majority of existing approaches (Gu et al., 2024; Hong et al., 2024), including recent works like Flow-GRPO (Liu et al., 2025) and DanceGRPO (Xue et al., 2025), treat the multi-step generation process as a “black box” with temporally agnostic optimization. They apply uniform updates across all timesteps despite the fact that each timestep operates under different noise conditions and contributes differently to final image quality. Specifically, we plot the Figure 2 (left) with applying SDE at only one timestep in the entire ODE trajectory, which ensures that any deviation in the final reward can be attributed to the stochastic exploration introduced at that specific step. As shown in Figure 2 (left), the std of the reward varies dramatically between timesteps, reaching a peak during early structural decisions (steps 0-2) and approaching zero during final refinements (steps 6-8). Yet Flow-GRPO maintains uniform treatment throughout, squandering high-impact exploration opportunities. Alternative approaches like SPO (Liang et al., 2025) attempt to address temporal dynamics through process reward models, but training such models on semantically ambiguous intermediate states is notoriously difficult. This raises a fundamental question: how can we effectively achieve precise credit assignment for intermediate actions while adapting optimization intensity to each timestep’s exploration capacity?
@@ 2 Related Work
Alignment for Diffusion Models. Alignment for diffusion models has emerged as a key research area. D3PO (Yang et al., 2024) introduces Direct Preference for Denoising Diffusion Policy Optimization to directly fine-tune diffusion models. Diffusion-DPO (Wallace et al., 2024) adapts Direct Preference Optimization (DPO (Rafailov et al., 2023)) as a simpler RLHF alternative that directly optimizes policies satisfying human preferences. DyMO (Xie & Gong, 2025) proposes a training-free alignment method for inference-time preference alignment. Flow-GRPO (Liu et al., 2025) first integrates online reinforcement learning into flow matching models. However, Flow-GRPO applies uniform optimization across timesteps and suffers from sparse terminal rewards, ignoring time-varying exploration potential in stochastic processes. Our TempFlow-GRPO addresses these limitations through trajectory branching for precise credit assignment and noise-aware policy weighting aligned with natural exploration capacity.
@@
Process Reward. Shaping reward processes beyond sparse terminal rewards significantly improves learning and policy performance. Zhang et al. (Zhang et al., 2025) combines response-level and step-level evaluation metrics. ThinkPRM (Khalifa et al., 2025) builds data-efficient process reward models (PRMs) using verification chain-of-thought for step-wise verification. In diffusion models, SPO (Liang et al., 2024) trains step-aware preference models for both noisy and clean images. However, PRMs require expensive step-level supervision. Methods like PRIME (Cui et al., 2025) enable online PRM updates using only outcome labels through implicit process rewards. Math-Shepherd (Wang et al., 2023) assigns reward scores to solution steps. Given challenges in scoring noisy images, algorithms enabling process rewards for flow models are needed. Our TempFlow-GRPO circumvents specialized process reward models by directly attributing outcome-based signals to intermediate actions, enabling precise credit assignment without computational overhead of training step-level evaluators for semantically ambiguous states.
@@ 3 Preliminary: Flow-GRPO
Given a prompt $\bm{c}$, the flow model $p_{\theta}$ samples a group of $G$ individual images $\{\bm{x}_{0}^{i}\}_{i=1}^{G}$ and the corresponding reverse-time trajectories $\{(\bm{x}_{T}^{i},\bm{x}_{T-1}^{i},...,\bm{x}_{0}^{i})\}_{i=1}^{G}$. Then, the advantage of the $i$-th image is calculated by normalizing the group-level rewards as follows:
@@
Figure 5: Overview of TempFlow-GRPO Framework. Our method performs trajectory branching by switching from ODE to SDE sampling at selected timesteps (t=k, j, i), injecting noise $\sigma_{t}\sqrt{\Delta t}\bm{\epsilon}$ to create exploratory branches. Each branch generates a distinct outcome with reward $R_{i}$, enabling precise credit assignment. The framework applies noise-aware weighting where $\omega_{i}>\omega_{j}>\omega_{k}$, prioritizing optimization at high-noise early stages (larger circles) over low-noise refinement phases (smaller circles), aligning learning intensity with each timestep’s intrinsic exploration capacity. We visualize the model’s learning process as an astronaut exploring unknown planets: in early stages , the model explores vast possibility spaces with high uncertainty, while later stages involve focused navigation toward the final destination.
@@ 4.1 Temporal Flow-GRPO
Flow-GRPO advances online RL for flow matching but overlooks time-dependent generative dynamics. We identify two key limitations: (1) Sparse Terminal Rewards: uniform credit assignment across timesteps fails to distinguish critical early decisions from later fine-tuning; (2) Uniform Optimization: ignoring non-uniform noise in SDE sampling, where high-noise early steps offer greater exploration potential than low-noise refinement phases (Xie & Gong, 2025). To address these issues, we propose TempFlow-GRPO (Figure 5), introducing process rewards via trajectory branching and noise-aware policy weighting.
@@ 4.1.2 Noise-Aware Policy Weighting
Reweighting by Noise Level. We wonder if the level of the exploration space itself can serve as a proxy for a reweighting factor. We visualized the noise level alongside the reward standard deviation (Figure 6, left) and observed a striking correspondence between the two. This strong correlation suggests that the noise level, serves as an excellent and intrinsic proxy for the exploration capacity and associated risk at each timestep.
We therefore propose to reweight the policy loss directly using the noise level. For each timestep $t$, we introduce a weighting factor proportional to the noise level. Specifically, we modify the original policy loss function to the following weighted form:
@@ 4.3 Group Strategy
Batch Group & Prompt Group. The original GRPO (Shao et al., 2024) framework employs a group-level strategy, where trajectories are grouped based on a shared prompt. To mitigate issues of overfitting and instability, Reinforce++ (Hu et al., 2025) later introduced a batch-level normalization.
Seed Group. In our work, trajectory branching requires $K$ distinct explorations at each timestep. Furthermore, we propose a novel seed-level grouping strategy. Under this approach, trajectories originating from the same prompt are further grouped if they share an identical initial noise. This methodology effectively controls for the influence of the initial noise, thereby ensuring that variations in reward can be attributed solely to the exploration conducted during the branching process. The experimental results, presented in Figure 7, validate our approach. As shown, TempFlow-GRPO consistently achieves superior performance compared to Flow-GRPO, regardless of the grouping strategy employed.
@@ 5 Experiment
Following Flow-GRPO, we validate our approach on the Compositional Image Generation benchmark in Geneval (Ghosh et al., 2023) and the Human Preference Alignment task in PickScore (Kirstain et al., 2023). To ensure a fair comparison, our experimental setup is kept consistent with that of Flow-GRPO. For instance, we normalized the weights applied to the policy loss to have a mean of 1 at all timesteps. While our method introduces a trajectory branching strategy, we maintain an identical group configuration by setting 4 initial noise seeds and a branching factor ($K$) of 6. This results in a total group size of 24 (4 seeds $\times$ 6 branches) and 48 groups, perfectly matching the setup in Flow-GRPO. Furthermore, to benchmark our method against concurrent work, we conduct additional experiments on the FLUX.1-dev model. More detailed experimental settings (include comparsion with DanceGRPO) can be found in the Appendix A.2.
@@ 5.1 Main Results
LDM (Rombach et al., 2022) - 0.37 0.92 0.29 0.23 0.70 0.02 0.05
SD1.5 (Rombach et al., 2022) - 0.43 0.97 0.38 0.35 0.76 0.04 0.06
SD2.1 (Rombach et al., 2022) - 0.50 0.98 0.51 0.44 0.85 0.07 0.17
SD-XL (Podell et al., 2023) - 0.55 0.98 0.74 0.39 0.85 0.15 0.23
DALLE-2 (Ramesh et al., 2022) - 0.52 0.94 0.66 0.49 0.77 0.10 0.19
DALLE-3 (Betker et al., 2023) - 0.67 0.96 0.87 0.47 0.83 0.43 0.45
@@
Show-o (Xie et al., 2024b) - 0.53 0.95 0.52 0.49 0.82 0.11 0.28
Emu3-Gen (Wang et al., 2024) - 0.54 0.98 0.71 0.34 0.81 0.17 0.21
JanusFlow (Ma et al., 2025a) - 0.63 0.97 0.59 0.45 0.83 0.53 0.42
Janus-Pro-7B (Chen et al., 2025) - 0.80 0.99 0.89 0.59 0.90 0.79 0.66
GPT-4o (Hurst et al., 2024) - 0.84 0.99 0.92 0.85 0.92 0.75 0.61
@@
FLUX.1 Dev (Black et al., 2025) - 0.66 0.98 0.81 0.74 0.79 0.22 0.45
SD3.5-L (Esser et al., 2024) - 0.71 0.98 0.89 0.73 0.83 0.34 0.47
SANA-1.5 4.8B (Xie et al., 2025) - 0.81 0.99 0.93 0.86 0.84 0.59 0.65
SD3.5-M (Esser et al., 2024) - 0.63 0.98 0.78 0.50 0.81 0.24 0.52
@@
SD3.5-M+Flow-GRPO (Liu et al., 2025) 5600 0.95 1.00 0.99 0.95 0.92 0.99 0.86
SD3.5-M+Flow-GRPO (Liu et al., 2025) 3800 0.88 0.99 0.96 0.90 0.87 0.83 0.78
@@
HPSv3 on FLUX.1-dev. To further validate the performance of TempFlow-GRPO across different model and reward models, we select FLUX.1-dev as our base model with 1024 resolution and employ HPSv3 (Ma et al., 2025b) as the reward model. As shown in Figure 8, which presents the results, our method requires only 80 steps to match the performance that Flow-GPO achieves in 300 steps. Moreover, our approach maintains a lower and more stable KL loss. Further experimental results and visualizations for the FLUX model can be found in the Appendix. A.3.
@@ 5.2 Analysis
Qualitative Result. We also conducted qualitative analyses on the FLUX.1-dev, Flow-GRPO (Prompt), and TempFlow-GRPO. As shown in Figure 10, compared to Flow-GRPO (Prompt), TempFlow-GRPO produces images with noticeably finer details and fewer visual artifacts or mistakes. In particular, our approach demonstrates superior capability in preserving complex structures and realistic textures. These qualitative improvements further highlight the advantages of our method in generating high-quality, visually appealing images.
@@ 6 Conclusion
Use of LLMs. We utilize LLMs to assist with formula derivations and writing refinement.
@@ A.1 Policy Gradient-Based Theoretical Framework
This reveals that the natural gradient coefficient is proportional to $\sqrt{\frac{1-k}{k}}\sqrt{\Delta k}$, which captures the intrinsic exploration potential at timestep $k$. After reweighting, we have the following derivation (don’t consider $a$ and norm coefficient):
@@ A.2 Experimental Setting Details
Dataset. We evaluate Compositional Image Generation on Geneval and Human Preference Alignment on PickScore. Meanwhile, we also test our method on HPDv2 (Wu et al., 2023).
Reward Model. In TempFlow-GRPO, we introduce four reward models, including clip-based PickScore, HPSv2 (Wu et al., 2023), VLM-based HPSv3 (Ma et al., 2025b) and framework Geneval.
Human Preference Alignment. Following Flow-GRPO, we evaluate our method on PickScore prompts using the PickScore reward function. PickScore is a CLIP-based scoring function that demonstrates superhuman performance in predicting human preferences for generated images. To ensure consistency with Flow-GRPO, we set the KL divergence regularization coefficient $\beta=0.001$.
Compositional Image Generation. We adopt the GenEval framework following Flow-GRPO’s experimental protocol. The training prompts are sourced from the Flow-GRPO dataset. GenEval provides an object-focused evaluation framework that assesses compositional image properties, including object co-occurrence, spatial positioning, object count, and color attributes. For fair comparison, we maintain the KL divergence regularization coefficient at $\beta=0.004$.